OpenStudy (princesssleelee):

Given that f(x) = x2 + 3x + 6 and g(x) = the quantity of x minus three, over two, solve for f(g(x)) when x = 1.

OpenStudy (kylexchaos):

4

OpenStudy (princesssleelee):

can you tell me how you got that please?

OpenStudy (kylexchaos):

Sure.

OpenStudy (kylexchaos):

x^2=(-1)^2=1 +3x=3(-1)=-3 =+6 adding these 1-3+6=4

OpenStudy (princesssleelee):

Thanks you

OpenStudy (fortytherapper):

Basically, lets say f(x) was 1+x and g(x) was 2+x, when x = 3. f(g(x)) basically says to plug g(x) where ever you see an x in f(x), so you would do 1+(2+x). You see it? Next, you would plug in 3 to find the answer, which would be 6. Same with your problem. f(x) is x^2+3x+6 and g(x) = (x-3)/2 So where you see a x in f(x), put the equation of g(x) y = ((x-3)/2)^2 + (3(x-3)/2) + 6 Which is completely bleh. You would then plug 1, which was x=1 the given, into that equation for your answer. y = ((1-3)/2)^2 + (3(1-3)/2) + 6 y = 1+(-3)+6 y = 4

OpenStudy (fortytherapper):

Late reply @princesssleelee My fault